# AttributionScanner: A Visual Analytics System for Model Validation with Metadata-Free Slice Finding

### v2

@@ Abstract
Model validation, visual analytics, data slicing, data-centric AI, human-assisted AI.
@@ I Introduction
As machine learning (ML) technologies continue to prevail in various domains, the importance of model interpretation and validation has grown significantly. Identifying a model’s failure and understanding the underlying reasons are crucial in promoting greater model accuracy, transparency, and accountability. A powerful technique supporting this effort is data slice finding [1], aiming to identify problematic data subgroups, referred to as slices, which often arise from under-represented samples or biased datasets [2]. These data slices are commonly defined by descriptive metadata, with contextual details of the model’s input such as feature descriptions, data collection details, and other relevant information.
After problematic data slices are identified, the remedial strategies often involve model training or fine-tuning with these slices to bolster the model’s overall reliability [3]. To enhance the trustworthiness, eXplainable Artificial Intelligence (XAI) [4] are also utilized in recent techniques to reveal potential causal factors adversely impacting the model and apply targeted model regularization [5] to improve the performance.
@@
Challenges of Data Slice Finding. State-of-the-art methods for data slice finding demand extensive metadata associated with each instance. Many works use pre-existing metadata collected together with the dataset [6, 7, 8], and others apply pre-trained vision-language models to generate descriptive data [3]. However, there are challenges hindering their applications—On the one hand, it requires significant effort to collect descriptive metadata. On the other hand, trained with general-purpose data, vision-language models can generate erroneous results in domain-specific scenarios. In situations where neither metadata nor suitable pre-trained models are available, obtaining interpretable data slices becomes challenging: requires manual examination of numerous instances or computationally expensive model training [9, 10].
Challenges of XAI for Vision Models. If troublesome data slices have been identified, it also requires non-trivial efforts to uncover the root causes of a mode’s failure. Even with the advanced human-friendly XAI techniques, there still exist a few challenges. Predominantly, most model explanation methods are instance-based, conducting a thorough examination of all instances to derive conclusions based on a big picture is practically unfeasible [11, 12, 13, 14]. Moreover, the insights we can obtain from existing methods are often not actionable, which means what the common failure patterns are and how to address them remain vague and demands further study.
Our approach. We present AttributionScanner, a Visual Analytics (VA) system designed to address the above challenges, providing an interpretable vision model validation framework with metadata-free slice finding. Our innovative workflow employs attribution-based explanation techniques to create interpretable feature representations for slice finding, sorting through unstructured image data into slices with common model behaviors. Moreover, we explain the discriminating features of each data slice via visual summaries, called Attribution Mosaic, to guide users in explaining model’s behaviors and detecting potential issues, thus facilitating model validation. Additionally, AttributionScanner integrates a human-in-the-loop approach to rank and annotate the problematic data slices based on common error types such as mislabeled samples and spurious correlations. Our approach enables users to apply these annotated slices to enhance the model, by correcting the identified errors on either the model or data side. Our main contributions include:
@@
  4. Qualitative and quantitative evaluation. We underscore the efficacy of AttributionScanner through two use cases with qualitative and quantitative evaluation, highlighting its ability to identify and mitigate hidden model issues.
This paper is structured as follows: in Sec. II, we review related work on data slice finding and XAI that is relevant to our work Sec. III outlines the methodology and system design of AttributionScanner. In Sec. IV, we illustrate two use cases showcasing how AttributionScanner can be utilized for vision model validation. An evaluation of our system and experts’ feedback are discussed in Sec. V. Lastly, we present our future work and conclusions in Sec. VII.
@@ II Related Work
Our method, AttributionScanner, leverages insights from both Data Slice Finding and XAI techniques to pinpoint problematic data slices in vision-based models. This provides a brief overview of the relevant literature on these two topics.
@@ II-A Data Slice Finding
Data slice finding technique is used to identify data subsets where ML models underperform. These subsets are usually defined by a combination of feature-value pairs extracted from tabular metadata. However, the exponential search space associated with this task makes it a challenging problem. To overcome this challenge, heuristics have been presented to approximate the solution. Barash et al. [15] introduced a Combinatorial Modeling-based approach that enumerates the space of value combinations in tabular data up to a set number of parameters. Similarly, Slice Finder [6] uses a decision tree heuristic to identify interpretable data slices that can be described by tree rules. Despite the advantages of heuristics, they may miss important data slices. Therefore, new techniques have been developed to enable the exhaustive search of data slices using novel formulations and parallel architectures. DivExplorer [1] employs frequent pattern mining to enumerate all candidate slices in the data and identify the most relevant ones based on frequency or metric divergence. In contrast, SliceLine [7] introduces a vectorized formulation for slice finding, which can be easily deployed in parallel devices.
The aforementioned methods necessitate structured metadata to segment datasets into meaningful subgroups, a requirement challenging to meet, especially in domains with large volumes of image or text data where detailed meta-information may be scarce. To address this, a prevalent approach involves encoding data samples into the model’s latent space, followed by clustering to group the data and pinpoint problematic slices. For instance, GEORGE [9] employs over-clustering alongside Group Distributionally Robust Optimization [16] to automatically detect and rectify distribution shift-induced errors. Spotlight [10] adopts a soft-clustering strategy, evaluating performance metrics per cluster and pinpointing clusters with the highest performance drops as data slices for user review. Similarly, UDIS [17] implements hierarchical clustering and further GradCAM [12] visual explanations for data slice examination. VisCUIT [18] allows a deeper exploration of the UDIS data slices by enabling slice comparison and the exploration of neuron activations per slice. Although these methods can support the identification of problematic slices, users have to delve into individual samples to understand model failure reasons. Domino [3] mitigates this by utilizing a pre-trained vision-language model for generating textual descriptions of slices, encoding images into model embedding, and clustering them based on shared visual attributes. However, this method’s reliance on pre-trained models for slice identification and explanation can backfire if there’s a significant domain mismatch between the dataset being evaluated and the pre-trained model, possibly yielding unmeaningful slices. In Sec. III, we introduce a methodology capable of producing explainable slices independently of such pre-trained models.
@@
Visualization has also proven effective in identifying problematic subgroups of data. Kahng et al. introduced a data cube analysis method [19], in which the performance of models was assessed on data slices defined by all pairwise combinations of data features. Meanwhile, FairVis [20] enables users to manually slice and inspect data slices based on domain knowledge to identify potential model biases. Similarly, Errudite [21] came up with a domain-specific language to allow the data slicing of textual documents based on linguistic features and model testing under specific conditions. Although manually slicing the data is useful for model analysis, this approach is not scalable since there could be an exponential number of slices to evaluate. To address this issue, SliceTeller [2] combined an automatic slice finding algorithm with a visual analytics tool to allow the analysis and iterative refinement of models. However, it still required structured metadata in order to produce meaningful data slices. In Sec. III, we present a visual analytics tool that requires no metadata or prior information to produce meaningful data slices.
@@
II-B XAI Techniques for Neural Networks
Given the widespread adoption of deep learning models, there is significant interest in understanding their prediction processes. In particular, visualization-based XAI methods have gained popularity due to their ability to produce intuitive and meaningful results and be applied to explain various deep Convolutional Neural Networks (CNNs). One genre is localization-based explanations, namely attribution heatmaps, which highlight image regions mostly contributing to a specified class prediction, providing class-discriminative localization to support model interpretation. Class Activation Maps (CAM) [11] is a pioneering method, but it requires a specific network architecture, namely the global average pooling layer, to generate visualizations. GradCAM [12] addresses this limitation by using gradients to compute model attributions. In our work, we utilize GradCAM attribution heatmaps in two ways: firstly, we use them to explain model predictions on single images. Secondly, we interpolate the attributions into weight matrices at the latent space to obtain attribution-weighted feature vectors, which are used to obtain and visualize data slices with similar model attribution patterns.
Another popular approach for understanding model behavior is optimization-based visual explanations. Feature Visualization [22, 13] creates images that maximally activate specific neurons, channels, or layers of a neural network, which focuses on understanding features that a model is sensitive to. Feature inversion [23] seeks to understand the information encapsulated in a group of spatial feature vectors. Specifically, it visually elucidates the shared patterns significant in those latent vectors from the model’s perspective.
@@ III-A Design Requirements
    Metadata-free. Our method should yield meaningful data slices without metadata, i.e., descriptive information other than the model’s training data.
@@
    Interpretability. We should provide sufficient and intuitive information to enhance model transparency. Diverging from prior metadata-free approaches that lack explanation capability, our system should enable both intuitive explanations of a data slice’s shared attributes and inspection of individual instances for insight verification as needed.
@@
    Actionable insights. The system should ensure the obtained insights are actionable, which requires careful design of visual guidance, issue annotation, and corresponding model regularization. The framework should allow for further expert engagement, such as re-labeling or model re-training, to enhance model performance.
@@ III-B System Workflow
AttributionScanner is a human-in-the-loop system, taking as input an image dataset and a trained CNN for image classification, and yielding interpretable data slices to assist human experts in metadata-free vision model validation.
@@ III-C System Interface
Figure 3: Attribution-weighted feature vector generation. An image is forwarded through CNN, where the corresponding feature vector $F$ and attributions heatmap $A$ can be extracted. Through linear interpolation of $A$ to the latent space, we obtain the weight matrix $W$ to calculate the attribution-weighted feature vector $F_{W}$.
@@ III-D Explainable Data Slice Finding
The process unfolds in three phases: initially, we create attribution-weighted feature vectors utilizing the interpolated attribution masks in the latent space. Subsequently, we derive an attribution representation space that maintains neighbor consistency in terms of model attributions and compute data slices. Lastly, we employ the Feature Inversion [23] to distill shared visual patterns per slice and visualize these patterns alongside model performance indicators in Attribution Mosaic.
@@
In this section, we illustrate the process of generating attribution-weighted feature vectors based on attribution heatmaps ( R1, R2). It is notable that AttributionScanner accommodates interchangeable attribution explanation techniques, such as CAM [11], GradCAM [12], or RISE [14] ( R2). Our prototype leverages GradCAM due to its demonstrated efficacy via sanity checks [26] and its prevalent usage for model validation in the ML community [5, 17, 27, 28, 29, 30, 31].
As shown in Fig. 3, an input image is processed by the model to extract the feature vector $F$, and a back-propagation is conducted from the model’s predictions to obtain the GradCAM attribution masks. While a conventional approach is to depict the mask as a heatmap for the instance-level explanation, we augment its explanation capability to the latent space to foster a more comprehensive understanding of the model’s attribution patterns. Specifically, we apply linear interpolation to map image-sized masks into the latent space, yielding a weight matrix $W\in\mathbb{R}^{m\times n}$. Then the normalized $W$ with $\Sigma_{i=1,j=1}^{i=m,j=n}W_{ij}=1$ is used to compute the weighted average of $F$. The resulting attribution-weighted feature vector $F_{W}\in\mathbb{R}^{1\times 1\times d}$ can be mathematically expressed as:
@@
To derive meaningful data slices exhibiting similar attribution patterns ( R1, R2), we construct an attribution representation space using the attribution-weighted feature vectors. Fig. 4 provides a qualitative comparison of the feature representation space (derived from the original feature vectors) and the attribution representation space (derived from the attribution-weighted feature vectors). For both spaces, we utilize UMAP [32] for dimensionality reduction to the 2D space. At each representation space, three points are randomly sampled and their nearest neighbors are obtained to examine the consistency of their corresponding attribution heatmaps.
As showcased in Fig. 4, it is obvious that the neighbor instances’ model attributions from the original space are not similar, while our constructed attribution representation space has greater local consistency concerning model attributions. Therefore, our attribution representation space provides a solid basis for computing interpretable data slices with coherent attribution patterns ( R2).
On top of our constructed attribution representation space, we proceed by clustering the samples with K-Means to obtain data slices. Similarly to DOMINO [3], we apply over-clustering by incrementing the number of clusters and monitoring the model attribution consistency within each slice, until a coherent grouping of samples is attained. This measure helps prevent the omission or overlooking of rare slices, ensuring a comprehensive slice identification.
@@ III-E Slice Summarization and Annotation
Although we have generated slices with consistent model attributions, validating the attribution correctness of each slice necessitates individual inspection of their encompassed data samples. This process is labor-intensive and time-consuming, posing the risk of overlooking or misidentifying critical issues. Therefore, a more efficient method to encapsulate the essence of each data slice is imperative. In response, we introduce Attribution Mosaic, a novel visualization technique that elevates Feature Inversion to visually summarize data slices in a mosaic depiction ( R3). We present the pipeline in Fig. 5. First, after the execution of data slice finding (Sec. III-D), we compute the convex hull of each slice based on the projected points in the 2D space. We then distill each slice’s predominant visual patterns via Feature Inversion, integrating the shape constraint of the convex hull. The final landscape featuring slices’ attribution summaries is presented as Attribution Mosaic.
@@
Slice Summary. A pivotal step in generating Attribution Mosaic involves visually summarizing the significant pattern in each slice ( R3). As discussed in Sec. II-B, Feature Inversion visualizes common patterns across a group of feature vectors [33, 34, 35]. We utilize this technique on the attribution-weighted feature vectors of each slice to visualize their shared attributions. Additionally, we enhance the quality of synthesized images by initializing them with the average of the original images instead of Gaussian noises [34] and parameterizing them in the Fourier space [35], making them better aligned with natural images rather than less-interpretable artificial patterns. The optimization process is defined by:
$x^{*}=\operatorname{arg\,max}_{x\in\mathbb{R}^{W\times H\times C}}\left(\ell(\phi(x),\phi_{0})+\lambda R(x)\right),$ (2)
@@
where $\phi:\mathbb{R}^{W\times H\times C}\rightarrow\mathbb{R}^{d}$ is the representation function; $\phi_{0}$ is a target feature activation; $\ell(*)$ is the loss function where the Euclidean Loss is used here, and $R(x)$ is a Total Variation (TV) regularization term to improve the quality of results [23]. In our scenario, we designate the optimization target, $\phi_{0}$, to be the average of attribution-weighted feature vectors of a data slice, and assign the representation function to the model’s convolutional modules. We obtain an image $x^{*}$ to minimize the loss function, thereby visualizing the shared attribution patterns within a data slice from the model’s perspective.
Mosaic Boundary Generation. The main purpose of this stage is to devise a clear layout enabling effortless exploration of diverse attribution patterns across slices. Merely obtaining and displaying the Feature Inversion results over the projected cluster centroids leads to overlapped visualizations, hindering effective user exploration. We design a mosaic layout, optimized for space-efficiency in information representation. Specifically, we compute convex hull [36] of each slice in the attribution representation space, yielding a convex hull for every slice with no overlap amongst them.
Mosaic drawing. Given that the convolutional modules take rectangular-shaped input, the original Feature Inversion algorithm solely generates rectangular visualizations. To preserve equivalent visualization in a mosaic form, we use our generated mosaic boundary as a constraint for optimization. Namely, we define the shape of $x^{*}$ as the minimal rectangle encompassing the corresponding convex hull, and in each iteration, we zero out pixels outside the hull, compelling the optimization to concentrate on the interior pixels. After rendering all feature visualizations, they are exhibited at their respective positions within the Attribution Mosaic.
@@
The Attribution Mosaic facilitates rapid comprehension of the main content of different slices, supporting the identification of potential errors in data or models ( R4). A prevalent and harmful issue in ML models, highly visible through Attribution Mosaic, is “spurious correlation”, where a model erroneously utilizes irrelevant image features for classification. For instance, in hair color classification, a model should make decisions according to hair features, termed as “correct/core correlation”, while the use of other features (e.g., background or facial attributes) denotes “spurious correlation.”
Such spurious correlations, regardless of model accuracy, pose substantial challenges including poor generalization in production [37] and AI fairness issues [38]. We introduce a metric termed Spuriousness probabilities to streamline the detection of spuriousness, ranging from 00 to $1$, indicative of the likelihood of spurious correlation existence. We employ the Label Spreading method [39] to propagate user annotated spurious/core information to other slices. Specifically, the minimal input requirement from users is one slice annotation of “core” (Spuriousness=00) or “spurious” (Spuriousness=$1$), and then AttributionScanner will propagate this information to create a Spuriousness Probability for the rest slices.

### v3

@@ Abstract
Model validation, data slicing, data-centric AI, human-assisted AI, visual analytics.
@@ I Introduction
As machine learning (ML) technologies continue to prevail in various domains, the importance of model interpretation and validation has grown significantly. Detecting a model’s failure and understanding the underlying reasons are crucial in promoting greater model accuracy, transparency, and accountability. A prevalent and harmful issue in ML models is “spurious correlation” [1], where a model erroneously utilizes irrelevant image features for classification. For instance, in hair color classification, a model should make decisions according to hair features, termed as “core correlation”, while the use of other features (e.g., background or facial attributes) denotes “spurious correlation.” Such spurious correlations, regardless of model accuracy, pose substantial challenges including poor generalization in production [2] and AI fairness issues [3].
However, models often perform inconsistently across different subsets of data, making it challenging to validate a model comprehensively to uncover hidden issues [4, 5, 6, 7, 8]. Thus, the crucial demands of identifying and characterizing such problematic data subsets have fueled data slice finding techniques [4, 5, 6, 7, 8, 9, 10]. The term “data slice” refers to data subgroups with coherent features, often defined by metadata, model representations, or other descriptive information. By detecting, understanding, and resolving issues in data slices, ML experts can improve model performance and ensure safer and more reliable deployments in high-stakes domains [11, 12, 13, 14]. To enhance the trustworthiness, eXplainable Artificial Intelligence (XAI) techniques [15] are also utilized to reveal potential factors adversely impacting the model and apply targeted model regularization [16].
@@
Challenges of Data Slice Finding. State-of-the-art methods for data slice finding demand extensive metadata associated with each single instance, which are often obtained from manual human annotations[17, 18, 19] or pre-trained vision-language models[14]. However, it is expensive to collect manually annotated metadata, and vision-language models can produce erroneous results, especially in domain-specific scenarios. When neither metadata nor suitable pre-trained models are available, obtaining data slices becomes challenging: requires manual examination of numerous instances or computationally expensive model training[20, 21].
Challenges of XAI for Vision Models. Explanation methods for vision models are often in instance-level[22, 23, 24, 25], producing explanations for individual data instances and leading to less-actionable insights—When observing a problematic model behavior, it is unknown whether it is a rare or common case, whether an actual model issue happens is still unclear. To obtain more reasonable conclusions, a thorough examination of all instances is required yet practically unfeasible.
Our approach. We present AttributionScanner, a Visual Analytics (VA) system designed for efficient vision model validation with slice finding, eliminating the costly requirements of metadata or other descriptive information. Our innovative workflow employs attribution-based explanation techniques to create interpretable feature representations for slice finding, enabling discriminative slice construction and intuitive data summary to streamline user exploration. Additionally, the methods provided by AttributionScanner are applicable to various computer vision models, collaboratively forming a human-in-the-loop approach for model validation and enhancement. Our main contributions include:
@@
This paper is structured as follows: in Sec. II, we review related work on data slice finding and XAI that is relevant to our work Sec. III outlines the methodology and system design of AttributionScanner. In Sec. IV, we illustrate two use cases showcasing how AttributionScanner can be utilized for vision model validation. An evaluation of our system and experts’ feedback are discussed in Sec. V. Lastly, we present our future work and conclusions in Sec. VI.
@@ II-A Data Slice Finding
Considering the fact the ML models often perform inconsistently across different data subsets, data slice finding technique is used to identify data subsets with coherent patterns, which can be defined by similar metadata, model representations, data statistics, or other descriptive information [26, 17, 9, 18]. By analyzing problematic data slices, ML experts can design targeted solutions to address model issues in edge cases, improving the model robustness and reliability more efficiently [4]. Existing approaches often necessitate expensive metadata to obtain meaningful subgroups, which hinders the applicability of data slice finding. To address this, a prevalent approach leverages feature vectors in the latent space and then clusters them to produce data slices [20, 21, 27, 28]. To provide data slice explanation, Domino [14] deploys a pre-trained vision-language model (VLM) to generate textual descriptions of slices, whose reliance on VLM can backfire if domain mismatch is significant, yielding less-meaningful slices. To fill the gap, we introduce a system providing explainable slices without metadata or pre-trained VLMs.
II-B Visual Analytics for Model Explanation and Validation
@@
Visual analytics have also proven effective for model explanation [29, 30, 31, 32, 33, 34, 35] and validation [36, 37, 38, 39, 40]. Using pairwise feature combinations, [41] produces data slices and assesses model performance with data cube analysis method. FairVis [42] enables users to manually slice and inspect data slices based on domain knowledge to identify potential model biases. Errudite [43] develops a domain-specific language for data slicing of textual documents. Although manually slicing the data is useful to a certain extent, it is not scalable because of its expansiveness. To address this issue, SliceTeller [10] combined an automatic slice finding algorithm with a visual analytics tool to allow the analysis and iterative refinement of models. However, it still required structured metadata to produce meaningful data slices. Our approach is for explainable model validation to eliminate the expensive metadata while involving effective human participation.
@@
II-C XAI Techniques for Neural Networks
Existing XAI techniques often use visualizations to support neural network interpretation, providing attribution heatmaps [22, 23, 24, 44] or visualizing models’ learned patterns [45, 46, 47]. Specifically, attribution-based explanations typically utilize information such as model gradients [23, 24], decision boundaries [22], or causal relationships [25, 44] to explain the model’s specific decision. However, despite the capability of offering intuitive and straightforward results, these methods are instance-based, explaining single data instances and missing the big picture of the common model behavior. On the other hand, methods such as Feature Visualization [45, 46] and Feature Inversion [47] create visualizations that maximally activate specific model neurons or layers, or visualizations that lead to feature vectors mostly similar to a target, which can be an aggregation of a group of feature vectors. In our approach, we leverage the benefit of Feature Inversion to highlight the shared patterns of a group of instances for visual summaries of our data slices, and also deploy GradCAM, a widely adopted attribution-based explanation method for explaining single instances.
@@ III-A Design Requirements
    Metadata-free. Our method should yield meaningful data slices without metadata, i.e., neither using metadata as input nor generating or extracting metadata.
@@
    Interpretability. We should provide sufficient and intuitive information to enhance model transparency, supporting interpretations of a data slice’s shared attributes for streamlined exploration and individual instances for insight verification as needed.
@@
    Actionable insights. The system should ensure the obtained insights are actionable, which requires careful design of visual guidance, issue annotation, and corresponding model or data regularization.
@@ III-B System Workflow
AttributionScanner is a human-in-the-loop system, taking as input an image dataset and a trained CNN, such as ResNet[48] for image classification, and yielding interpretable data slices to assist experts in efficient model validation.
@@ III-C System Interface
Figure 3: Attribution-weighted feature vector generation. An image is forwarded through CNN, where the corresponding feature vector $F$ and weight matrix $W$ can be extracted. Through linear interpolation of $W$ to the image space, we obtain the attribution heatmap $A$ to calculate the attribution-weighted feature vector $F_{W}$.
@@ III-D Explainable Data Slice Finding
The process unfolds in three phases: initially, we create attribution-weighted feature vectors utilizing attribution masks in the latent space. Subsequently, we derive an attribution representation space and compute data slices. Lastly, we employ the Feature Inversion [47] to distill shared visual patterns per slice and visualize these patterns alongside model performance indicators in Attribution Mosaic.
@@
In this section, we detail the generation of attribution-weighted feature vectors ( R1, R2). It is notable that there are existing approaches employing attribution-weighted rather than original feature vectors, which aim to provide better model explanation [47, 50, 51] or conduct explainable model enhancement [16, 7]. Considering attribution-weighted feature vectors reserve both data features and model attributes [51, 16, 7], we opt for this design choice to support our explainable data slice construction. Specifically, we leverages GradCAM to obtain model attributes in the latent space, due to its demonstrated efficacy via sanity checks [52] and prevalent usage [16, 27, 53, 54, 55, 56, 57]. In our approach, GradCAM is interchangeable with other XAI techniques such as CAM [23] or RISE [25] ( R2).
As shown in Fig. 3, an input image is processed by the model to extract the feature vector $F$, and GradCAMis performed to produce attributes in the latent space $W$, which can be linearly interpolated to the image size for attribution heatmap. $W\in\mathbb{R}^{m\times n}$ and is normalized with $\Sigma_{i=1,j=1}^{i=m,j=n}W_{ij}=1$ to compute the weighted average of $F$. The resulting attribution-weighted feature vector $F_{W}\in\mathbb{R}^{1\times 1\times d}$ is:
@@
Representation Space Construction. To derive meaningful data slices ( R1, R2), we need to construct a representation space with neighbor consistency in terms of model attributes and data features. We first conduct quantitative and qualitative experiments to validate whether attribution-weighted feature vectors help fulfill our requirement. Specifically, we compare the feature representation space (derived from the original feature vectors) and the attribution representation space (derived from the attribution-weighted feature vectors) As shown in Fig. 4, following the established practices[14, 4, 5] and considering the time efficiency [58, 59], we utilize dimensionality reduction by UMAP [58] to produce both spaces. At each space, three points are randomly sampled and their nearest neighbors are obtained. We qualitatively examine the consistency of their corresponding attribution heatmaps. As showcased in Fig. 4, neighboring instances’ model attributions on the original space are not similar, while our constructed attribution representation space has greater local consistency concerning model attributions.
In addition, we quantitatively compute the neighboring consistency regarding data features and model attributions and report the results in Fig. 4. For each point, we compute the average cosine similarities between its top 10 neighbors using their original feature vectors, representing data feature consistency; besides, we compute the average Wasserstein similarities of their attribution mask, where higher values of both scores indicate better consistency. Lastly, we compute the average scores to quantify the overall neighbor consistency of data features and model attributions for each space. As shown in Fig. 4, compared to the original feature space, our constructed space has comparable data feature similarity (0.8987) and significantly higher attribution similarity (1.0659), which proves that the attribution-weighted feature vectors help construct a space with satisfactory local consistency in terms of data features and model attributions, providing a solid basis for computing interpretable data slices ( R2).
Data Slice Identification. On top of the attribution representation space, we apply K-Means clustering to obtain data slices. Similarly to DOMINO [14], we apply over-clustering by incrementing the number of clusters and monitoring the model attribution consistency within each slice, until a coherent grouping of samples is attained. Specifically, the attribution consistency of a data slice is measured as the average cosine similarity of each instance’s attribution-weighted feature vector with the slice centroid. We set a threshold at $0.8$ to determine slice coherence in our case studies. Note that we allow ML experts to adjust this threshold based on the specific requirements of their datasets and tasks. This helps prevent the omission or overlooking of rare slices and provides a comprehensive slice identification.
@@ III-E Slice Summarization and Annotation
After obtaining slices with consistent model attributions, we design Attribution Mosaic, a novel visualization technique, to visually summarize data slices in a mosaic depiction( R3). We present the pipeline in Fig. 5. First, we compute the convex hull of each slice based on the projected points in the 2D space. We then distill each slice’s predominant visual patterns via Feature Inversion, integrating the shape constraint of the convex hull. The final landscape featuring slices’ attribution summaries is presented as Attribution Mosaic.
Mosaic Boundary Generation. Given that the convolutional modules take rectangular-shaped input, conventional Feature Inversion produces rectangular visualizations, merely obtaining and displaying its results over the projected cluster centroids leads to overlapped visualizations, hindering effective user exploration. Since K-Means is a centroid-based algorithm that partitions the space into strict non-overlapping Voronoi cells, this guarantees that the convex hulls [60] of our data slices have no overlap.
Therefore, we compute the convex hull of each slice on the attribution representation space, which provides a mosaic layout for our slice visualization.
@@
Mosaic Drawing with Feature Inversion. We deploy Feature Inversion technique [61, 62, 63] to visualize the significant pattern in each slice ( R3), which is an optimization-based technique to visualize shared features across a group of feature vectors. Specifically, we first create an image $I$ with size $W\times H$, which is the minimal rectangle encompassing the mosaic boundary of a specific slice denoted by $M$. The optimization process focuses on updating pixels in $I$ that lie inside $M$, synthesizing a mosaic-shaped image to visualize the dominant pattern of the feature vectors. To enhance the quality of synthesized images, we initialize them with the average of the original images instead of Gaussian noises [62] and parameterize them in the Fourier space [63], making them better aligned with natural images rather than less-interpretable artificial patterns. The optimization process is defined by:
$x^{*}=\operatorname{arg\,max}_{x\in M}\left(\ell(\phi(x),\phi_{0})+\lambda R(x)\right),$ (2)
@@
where $x^{*}$ denotes the pixels in $I$ bounded by the mosaic shape $M$, the convex hull of the slice. The values of $x^{*}$ are updated iteratively through the optimization process to maximize the target function. $\phi:\mathbb{R}^{W\times H\times C}\rightarrow\mathbb{R}^{d}$ is the representation function, i.e., the model’s convolutional modules. $\phi_{0}$ is a target feature activation, which is the average of attribution-weighted feature vectors of a data slice in our case. $\ell(*)$ is the loss function where the Euclidean Loss is used here, and $R(x)$ is a Total Variation (TV) regularization term to improve the quality of results [47]. Thus, we obtain a mosaic-shaped image to visualize the shared attribution patterns within a data slice. Lastly, the visualizations from all data slices are exhibited at their respective positions within the Attribution Mosaic.
@@
The Attribution Mosaic facilitates rapid comprehension of the main content of different slices, supporting the identification of potential spurious correlation issues( R4).
We introduce a metric termed Spuriousness probabilities to streamline the detection of spuriousness, ranging from 00 to $1$, indicative of the likelihood of spurious correlation existence. We employ the Label Spreading algorithm [64] to propagate user-annotated spurious/core information to other slices. Specifically, a graph is constructed where nodes represent data points and edge weights represent similarities between points. In our case, we fit the graph to our 2D attribution representation space.
@@
According to the graph, the Label Spreading algorithm updates the labels of unannotated slices iteratively to ensure local and global label consistency—similar slices should have similar spuriousness, and the propagated spuriousness scores should respect the annotated slice spurious/core labels. The minimal input requirement from users is one slice annotation of “core” or “spurious” (Spuriousness=00 or $1$), and then AttributionScanner will propagate this information to create Spuriousness probabilities for the rest of slices.